OpenStudy (anonymous):

Lily is practicing multiplying complex numbers using the complex number (2 + i). To determine the value of (2+i)^2 , Lily performs the following operations. Lily made an error. Explain Lily's error and correct the step which contains the error.

OpenStudy (anonymous):



OpenStudy (phi):

step 1 says \[ (2+i)^2 = 4+i^2\] that is wrong.

OpenStudy (phi):

\[ (2+i)^2 \ \text{ means } (2+i)(2+i) \] I would multiply that out the same way you would multiply for example (a+b)(a+b)

OpenStudy (anonymous):

so that part just cancels itself out?

OpenStudy (phi):

do you know how to multiply (x+y)(x+y) ? some people use FOIL or use the distributive property for example, let A be the first (x+y). we (should!) know A(x+y) can we written as A*x + A*y but A is short for (x+y) so that is (x+y)x + (x+y)y now distribute again: x*x + xy + xy + y^2 or x^2 +2xy + y^2 we use that same idea on (2+i)(2+i)

OpenStudy (anonymous):

okay so the first step would then look like (2+i)^2 = 2^2+i^2+4i

OpenStudy (phi):

yes, exactly next, use the simplification that i*i is -1 (by definition) so you get \[ (2+i)^2 = 2^2+i^2+4i \\ = 4-1+4i \\ = 3+4i \]

OpenStudy (anonymous):

okay so is that all the step redone or is there more that needs to be changed?

OpenStudy (phi):

the rest of the steps up above are ok, but don't make sense because they started wrong. I would say: step 1: (2+i)(2+i) = 4 + 4i+i^2 step 2: i^2 =1 so 4 + 4i-1 step 3: 3+4i

OpenStudy (anonymous):

okay thank you so much!